Recently, the Transport Association of Zambia increased the price of bus fares for all long routes in Zambia. Specifically, the price of tickets between Lusaka and Livingstone increased from 200 Kwacha to 250 kwacha. The demand equation Q_d = 2000 - 5P is representative of the demand for bus tickets between Lusaka and Livingstone.

(i)Calculate the new price quantity demanded level
1
Expert's answer
2021-09-07T09:46:43-0400

Given

new price 250

"Qs=2000-5p"

Then

"Q=2000-5(250)\\=2000-1250\\=750"